1. Nitrogen and phosphorus (N and P) are important elements found in fertilizer that encourage health plant growth. As part of an experiment, a student requires some fertilizer for their plants. The student can choose between a 10.0 kg bag of KNO3 or a 12.0 kg bag of LiPO2. Assume N and P help plants grow equally well. 


Calculate the mass % of N and P.


Determine which bag of fertilizer, KNO3 or LiPO2 will support greater plant growth.


1
Expert's answer
2021-04-27T07:54:38-0400

For KNO3

MM = 102.1032g/mol


Atomic mass of N = 14.0067 g/mol


% by mass = "\dfrac{14.0067}{101.1032}x100 = 13.87" %


Mass of N in 10kg of KNO3 = "\dfrac{14.0067}{101.1032}x10kg = 1.385kg"



For LiPO2

MM = 69.9136g/mol


Atomic mass of P = 30.973762g/mol


% by mass = "\dfrac{30.973762}{69.9136}x100 = 44.30" %


Mass of P in 12kg of LiPO2 = "\dfrac{30.973762}{69.9136}x12kg = 5.316kg"


Conclusion 12kg of LiPO2 contains more of the important nutrient hence will support more growth than 10kg of KNO3